Pixel:

Your medical bill is $2345. Your health insurance covers 70% after a $120 deductible. What amount of the bill will you pay?

Pixel:

2345 - 120 = 2225 2225 / .7 = 1557.50 2345 - 1557.50 = 787.50 787.50 + 120 = 907.50

Pixel:

@TheSmartOne

Pixel:

i think i did this wrong

Pixel:

wait i already added in the deductible, 787.50 is the answer.?

TheSmartOne:

No, you did it correct the first time You accidentally put a / instead of * but you multiplied which is correct. I think you're still having some tricks so I'll try to explain it again You have this bill Your have to pay your deductible first Insurance will then pay their share, your pay the remaining The total amount you pay can be calculated a few different ways You found how much the insurance is paying after the deductible. Then you subtract and find out how much you were paying. Which is what you did. It would be much easier to just multiply by how much you have to pay which is going to be 100% - the percentage insurance is posting But anyways then you add in the deductible because that amount we just calculated only consists of how much you pay after you paid the deductible

TheSmartOne:

Does that make more sense

Pixel:

i guess, i still got the answer in the end

TheSmartOne:

Ye just trying to save you some time and work